OpenStudy (anonymous):

2 = 2t – 2 i will give medal+fan

OpenStudy (anonymous):

$$2 = 2t – 2$$ $$2 + 2 = 2t$$ $$4=2t$$ Divide both sides by two: $$\frac{4}{2}=\frac{2t}{2}$$ $$2 = t$$ Check it: $$2 = 2(2) – 2$$ $$2 = 4 - 2$$ $$2 = 2$$ $2 = 2$ = True!